OpenStudy (anonymous):

Use symbols to write the expression and then evaluate it thirty-six squared divided by six cubed

OpenStudy (compassionate):

\[36^2\div6^3\]

OpenStudy (anonymous):

Thank You

OpenStudy (zehanz):

To evaluate, remember that 36 is 6², so you have to calculate\[\frac{ 36^2 }{ 6^3 }=\frac{ (6^2)^2 }{ 6^3 }=...\]Well, you can see where this leads to... ;)

OpenStudy (zehanz):

Just promise not to use a calculator!!

OpenStudy (anonymous):

= 6

OpenStudy (zehanz):

Well done!